OpenStudy (anonymous):

Given the following function, which statements are true? F(x)={-x+1,<0} {-2,x=0} {x^2-1>0} A) f(4)=7 B) f(-2)=0 C)f(1)=0 D)f(-1)=2

OpenStudy (jango_in_dtown):

f(1)=1^2-1=0 so option C is correct

OpenStudy (jango_in_dtown):

Als f(-1)=-(-1)+1=2 so option D is also correct